Computer Warehouse sells batteries ($3) and small boxes of pens ($5). In July, total sales were $960.
Customers bought 5 times as many batteries as boxes of pens. How many of each did Computer Warehouse sell?
Answer by stanbon(75887) 
 Computer Warehouse sells batteries ($3) and small boxes of pens ($5). In July, total sales were $960.
Customers bought 5 times as many batteries as boxes of pens. How many of each did Computer Warehouse sell?
--------------------
Quantity equation: b = 5p
Value Equation: 3b + 5p = 960
---------------------------------
Substitute to solve for "p":
3(5p) + 5p = 960
15p + 5p = 960
20p = 960
p = 48 (# of boxes of pens sold)
-----
Since b = 5p, b = 5*48 = 240 (# of batteries sold)
======================================================
Cheers,
Stan H.
|
|
|